# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v2

@@ 4.3 Welfare analysis
We follow Krusell et al. (2010) in measuring the welfare change for black and white workers by calculating the change in average consumption equivalence after removing discriminatory hiring. [^11] Under the benchmark model, let $V(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(c_{t})$ be the maximal value of the individual with employment status $e$, race $R$, productivity $s$, and asset $a$. Under an alternate economy, let $\tilde{V}(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$ be the maximal value of individuals with each corresponding state. We examine the welfare change between the two economies through consumption equivalence $\Omega$, following the equation: $\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log((1+\Omega)c_{t})=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$. With $\log$ utility, we derive $\Omega=\exp((\tilde{V}-V)(1-\beta))-1$. We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change. We aggregate over the counterfactual distribution rather than the benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers associated with the states of working at and unemployment from $p$ firms. Aggregating over the benchmark distribution overstates the welfare change for white workers without accounting for the distribution shift. We present the average welfare changes for black and white workers by productivity types and wealth quintiles in Table 5.
Overall, the economy experiences a 5.83% reduction in average welfare if we eliminate $p$ firms and a 1.17% increase in average welfare if we separate the job destruction shocks by race. Once we remove discriminatory hiring, black workers experience an average welfare increase of between 6.89% and 10.94%. This is because of the rise in their job opportunities and higher wage rates. On the other hand, white workers encounter an average welfare loss of between 1.14% and 8.83%. This is because of the decrease in their wage rates and the increase in the unemployment rate.
Mid-panels in Table 5 present the heterogeneous welfare change for black and white workers by productivity types. Though the welfare reduction from eliminating prejudiced firms is similar across productivity for white workers, the most considerable loss is for the lowest productive white workers at 9.03% when we remove $p$ firms. In comparison, the highest productive workers lose the most at 1.19% when further adjusting the job destruction rate by race. Furthermore, the most productive black workers have the highest welfare gain at 7.41% and 11.49% across the two counterfactual models.
@@
Average welfare gain (%) Eliminate $p$ firms Race specific $\lambda$
Overall -5.83 1.17
Black White Black White
6.89 -8.83 10.94 -1.14
@@
Low 20% 6.85 -9.43 10.98 -1.13
40-60% 6.95 -8.95 10.94 -1.17
Top 20% 6.88 -8.14 10.72 -1.09
Note: This table compares the average consumption equivalence change from the benchmark steady state to the equilibrium where we eliminate $p$ firms (Columns 1 and 2) and further adjust the job destruction shock to be race-specific (Columns 3 and 4). Low, mid, and high productivity corresponds to the lowest, middle, and highest value of idiosyncratic productivity $s$. The wealth quintiles are based on benchmark steady-state wealth distribution.
The bottom panels of Table 5 present the welfare change by wealth quintiles for black and white workers. Removing prejudiced firms benefits the mid-quintile black workers the most, at 6.95%, and the lowest-quintile black workers the most, at 10.98%, after adjusting the job destruction shocks by race. The most considerable reduction in welfare (9.43%) occurs for the lowest quintile white workers in the model removing $p$ firms and for mid-quintile white workers at 1.17% after adjusting for race-specific job destruction shocks.
@@ 5 Business cycle dynamics
In this section, we begin by documenting the racial disparities over business cycles. Next, we incorporate aggregate uncertainty into our benchmark model to quantify the impact of hiring discrimination on generating racially disparate dynamics within these cycles. Finally, we examine the welfare consequences of hiring discrimination during business cycles.
@@ 5.1 Racial disparities over the business cycles
Table 6 presents the business cycle disparities in labor market outcomes between black and white workers. The top panel shows the dynamics of the racial unemployment rate gap, and the bottom panel shows the dynamics of the average wage. Cyclicality is measured by the correlation between the variable and real GDP. Volatility is calculated by the standard deviation of the variable relative to the standard deviation of real GDP.
Unemployment rate: We use monthly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the black-to-white unemployment rate difference. We then log the values and apply the HP filter with a smoothing parameter of 1600 to extract cyclical components. We find that the racial unemployment rate gap is countercyclical, widening during recessions and narrowing during expansions. It has a relative volatility of 9.71.
@@
Black-White gap -0.6255 9.7100
@@
Black 2.6041 2.5352
White 1.2729 2.0848
Note: This table provides business cycle statistics for the US. The unemployment rate was obtained from the Current Population Survey (CPS) from 1996 to 2014. Average wage statistics are constructed from the Panel Study of Income Dynamics (PSID) of the same period. All variables are logged and HP filtered with a smoothing parameter of 1600.
Wage: We rely on the individual data in PSID from 1996 to 2014 to construct a wage panel to calculate the cyclical properties of the average wage. This is due to the composition change of workers over the business cycles. Using aggregate wage data underestimates its cyclical properties (Stockman, 1983; Bils, 1985; Solon et al., 1994). For the same reason, we report the cyclical properties of average wages for black and white workers separately rather than calculating the cyclical properties for a wage differential index. Wage cyclicality is calculated through a wage regression as the percentage change of average wage when real GDP increases by one percent after controlling for demographics. Wage volatility is calculated by the relative standard deviations of the year-to-year change in wages from the wage regression. [^12] We conduct wage regression by regressing the change in individual log wage on the change in log GDP, after controlling for demographic factors. Specifically, we follow Bils (1985), Solon et al. (1994), and Devereux (2001) and run the regression: $\Delta\log w_{it}=\alpha+\beta\Delta\log Y_{t}+\gamma x_{it}+\nu_{t}+\upsilon_{s}+\epsilon_{it}$ (22) where $Y_{t}$ refers to real GDP. Demographic variables $x_{it}$ include gender, education, labor market experience, and marital status. We also include year fixed effects, $\nu_{t}$, and state fixed effects, $\upsilon_{s}$. We measure cyclicality using coefficient $\beta$ from the regression. The volatility is the standard deviation of the year fixed effects, $\nu_{t}$. We then adjust the estimation to quarterly frequency. We find that wages for black workers are twice as cyclical as wages for white workers once we control for demographics. Black workers’ average wage is more volatile (2.54) than that of white workers (2.08).
@@ 5.2 Augmented model with aggregate shocks
We introduce an aggregate total factor productivity (TFP) shock $z$ to the benchmark steady-state model from Section 2. The TFP shocks follow an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$. The output of the matched firm-worker pairs depends on the realizations of aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
With the introduction of aggregate shocks, the state space expands to include aggregate states $(z,\mu)$, where $\mu$ is the distribution of workers across employment status ($e$), race ($R$), idiosyncratic productivity ($s$), and asset ($a$). The next period’s aggregate distribution $\mu^{\prime}$ is determined by $(z,\mu)$, and the law of motion is given by $\mu^{\prime}=\Gamma(z,\mu)$.
We follow Cooley et al. (1995) and Boppart et al. (2018) and choose the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015. All other parameters remain the same as in Table 1. This gives us an aggregate output volatility of 0.02 and a first-order auto-correlation of 0.73.
We obtain the stochastic equilibrium of our model by using the sequence space method of Boppart et al. (2018). We first solve non-linearly for the perfect foresight transitions to a single small MIT shock, i.e., an unexpected shock to the aggregate TFP. We then use the solved impulse responses as the numerical derivatives with respect to the initial TFP shock. Using these derivatives, we simulate the stochastic equilibrium by generating the TFP realizations and calculating the corresponding model moments over the business cycles as a linear combination of the impulse response and the TFP realizations. [^13] This method hinges on the assumption that the business cycle dynamics can be well approximated as a linear system. In Appendix C, we demonstrate the validity of this assumption by establishing the symmetry of the impulse responses to 1% positive and negative TFP shocks.
@@
Black-White gap -0.6255 -0.6731 9.7100 0.3009
@@
Black 2.6041 1.0749 2.5352 1.0755
White 1.2729 0.9984 2.0848 0.9985
Note: This table compares the data moments and the benchmark generated moments on the black-white unemployment and wage gaps over the business cycles.
Table 7 compares data and model moments on unemployment and wages. Similar to standard search models, the stochastic dynamics of our model are subject to the Shimer Puzzle (Shimer, 2005). The augmented benchmark model generates a similar cyclicality in the unemployment rate gap as the data but with much less volatility in the unemployment rate gap. Regarding the average wage, the model has lower cyclicality for both black and white workers and less pronounced differences in volatility. However, qualitatively, we observe a higher wage cyclicality and volatility for black workers than for white workers.
@@ 5.3 Impact of hiring discrimination over business cycle
We compare the benchmark dynamics with the counterfactual model without $p$ firms to examine the role of discriminatory hiring in the context of aggregate fluctuations. [^14] As with Table 4, the counterfactual economy is a one-sector economy with only $np$ firms. We do not further adjust race-specific destruction shocks.
@@ 5.3.1 Impulse responses
Figure 3 illustrates the labor market, wealth, and consumption responses of black and white workers following a 1% unexpected expansionary TFP shock. We evaluate the benchmark model’s reaction against the counterfactual model that excludes $p$ firms. In general, our models produce countercyclical unemployment rates alongside procyclical patterns in wages, wealth, and consumption.
@@
Note: This figure plots the impulse response functions of unemployment rates, black-white average wage and wealth ratios, and aggregate consumption between black and white workers to a one percent increase in aggregate TFP at date zero. The red solid line denotes the responses from the benchmark model, and the blue dashed line denotes the responses from the counterfactual model without $p$ firms.
In the model without $p$ firms, the unemployment rate of black workers decreases less than that of the benchmark economy after the expansionary shock. This implies that their unemployment rate is less volatile without $p$ firms. This is because $np$ firms have to compete against $p$ firms for hiring, so they post more vacancies, providing more opportunities for all workers. For white workers, the difference between the benchmark and the model without $p$ firms is much smaller. Interestingly, their recovery in the counterfactual model is much faster than in the benchmark model. [^15] This is a common property in the literature on dual labor markets. A single-market model often reverts to its steady state faster than a dual-market model (e.g. Horvath and Yang, 2022). This is because the $p$ firms have a low job destruction shock. Once hired, white workers are less likely to lose their jobs, hence a slower rate to a steady state.
The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than without prejudiced firms. This suggests more volatile procyclical wage and wealth responses for black workers than for white workers in the presence of $p$ firms. Lastly, we don’t observe a discerning difference in the consumption responses with and without $p$ firms.
@@ 5.3.2 Stochastic simulation
We simulate the stochastic equilibrium of the economy over 10,000 periods under aggregate TFP shocks, comparing outcomes with and without $p$ firms. While impulse responses to an MIT shock capture the transition path following a one-time disturbance, the stochastic equilibrium focuses on the economy’s short-term responses to continuous aggregate shocks. The results are summarized in Table 8.
In a counterfactual equilibrium without $p$ firms, the racial unemployment rate gap closes, eliminating its cyclical characteristics. Likewise, the disparities in the cyclicality and volatility of average wages between racial groups are significantly diminished. According to Table 8, the benchmark model reveals a pronounced countercyclicality (-0.6731) and volatility (0.3009) for the black-white unemployment rate gap, both of which disappear when $p$ firms are not included. For average wages, the cyclicality for black workers lowers from 1.0749 in the benchmark to 1.0131 without $p$ firms, whereas it increases slightly for white workers from 0.9984 to 1.0115. Conversely, the volatility for black workers drops from 1.0755 to 1.0139, while for white workers, it rises from 0.9985 to 1.0116. This convergence indicates that hiring discrimination significantly contributes to the cyclical disparities in the labor market between black and white workers. These results are consistent with Cajner et al. (2017), Ganong et al. (2020), and Ragusett (2022), which argue that demographic differences alone do not sufficiently explain the racial disparities observed throughout the business cycle.
Table 8: Impact of hiring discrimination on business cycle racial gaps
Cyclicality Volatility
@@
Black-White gap -0.6731 0.0000 0.3009 0.0000
@@
Black 1.0749 1.0131 1.0755 1.0139
White 0.9984 1.0115 0.9985 1.0116
Wealth ratio
Wealth ratio 0.2028 0.2054 0.1611 0.1466
Consumption ratio
Consumption ratio 0.9717 0.9695 0.1861 0.1659
Note: The top panel compares the black-white unemployment rates, wages, consumption, and wealth gaps over the business cycles between the benchmark model and the model without $p$ firms.
We further examine the cyclical properties of the average black-to-white wealth ratio and consumption ratio. In the benchmark model, the cyclicality of the average wealth ratio is 0.2028, which changes slightly to 0.2054 without $p$ firms, while its volatility decreases from 0.1611 to 0.1466. For the average consumption ratio, the cyclicality decreases marginally from 0.9717 to 0.9695, and the volatility drops more significantly from 0.1861 to 0.1659. These results show that, aside from the cyclicality of the wealth ratio, the cyclical properties of racial gaps across all other dimensions are notably reduced when discriminatory hiring is eliminated.
In summary, $p$ firms drive racial disparities in the labor market over the business cycle. The black-white unemployment rate gap, countercyclical in the benchmark, vanishes without $p$ firms. Wage cyclicality and volatility for black workers also decline, converging toward those of white workers. While the wealth ratio’s cyclicality rises slightly, its volatility and that of the consumption ratio decrease. These findings indicate that hiring discrimination is a key factor shaping racial differences in the cyclical properties of the labor market and economic outcomes.
5.4 Welfare disparities in business cycles
Lastly, we examine the business cycle welfare implications of removing hiring discrimination. We simulate the economy for 10,000 periods and compute the long-run average welfare for each individual type, following Cho and Ma (2023). The welfare measure is calculated as average consumption equivalence described in Section 4.3 over the simulation periods. We present the results in Table 9.
Table 9: Heterogeneous welfare change in business cycles
Average welfare gain (%) Eliminate $p$ firms
Overall -0.01
Black White
Average 0.02 -0.01
by productivity
Low 0.00 -0.01
Mid 0.02 -0.01
High 0.02 -0.02
by wealth
Low 20% 0.01 -0.01
40-60% 0.02 -0.02
Top 20% 0.03 -0.02
Note: This table compares the average consumption equivalence change from the benchmark model to the one without $p$ firms in business cycles. We simulate the economy for 10,000 periods and calculate the long-run average welfare for each individual type. Low, mid, and high productivity correspond to the lowest, middle, and highest values of idiosyncratic productivity $s$. The wealth quintiles are based on benchmark steady-state wealth distribution.
The economy experiences an average welfare decrease of 0.01% after removing prejudiced firms. This is much less than the welfare change in the steady state in Table 5, primarily because workers are risk averse. All workers experience a decrease in unemployment rate volatility, while white workers experience an increase in wage volatility that somewhat counters the changes in other dimensions. On average, black workers experience a modest welfare gain of 0.02%, whereas white workers face a slight welfare loss of 0.01%. Among white workers, those with high productivity and those in the top 20% of the wealth distribution lose the most (0.02%). In contrast, black workers in the top wealth quintile benefit the most, with an average welfare gain of 0.03%.

### v3

@@ 4.3 Welfare analysis
We follow Krusell et al. (2010) in measuring the welfare effects of discriminatory hiring on black and white workers. To do so, we compute the change in average consumption equivalence after removing $p$ firms. [^12] See Appendix C for details on the welfare calculation. Table 5 presents the overall welfare change, as well as the breakdown by productivity type and wealth quintile for both groups.
Eliminating hiring discrimination raises aggregate welfare by 1.5%. Black workers experience a 12.9% increase in their average welfare, mainly due to improved job opportunities and higher wage rates. White workers, on the other hand, experience a modest 0.48% decline, driven by a slight drop in wages and a higher unemployment rate after losing access to the exclusive $p$ firms. Although white workers make up about 85% of the population, the substantial gains for black workers more than offset the small loss among white workers, resulting in a net rise in overall welfare.
@@
Average welfare gain (%) No $p$ firms
Aggregate 1.51
Black White
Overall 12.89 -0.48
@@
Low 20% 12.91 -0.45
40-60% 12.92 -0.49
Top 20% 12.65 -0.48
Note: This table compares the average consumption equivalence change from the benchmark steady state to an equilibrium with no $p$ firms. Low, mid, and high productivity corresponds to the lowest, middle, and highest value of idiosyncratic productivity $s$. The wealth quintiles are based on benchmark steady-state wealth distribution.
The middle panel in Table 5 displays the heterogeneous welfare changes for black and white workers by productivity. For white workers, the welfare decline from eliminating prejudiced firms is similar across productivity levels, with a slightly larger loss for those in the middle productivity group. In contrast, among black workers, the highest productivity group experiences the greatest gain (13.57%), followed by the middle and low productivity groups. The bottom panel in Table 5 breaks down the welfare changes by wealth quintile. Removing prejudiced firms benefits black workers most in the middle quintile, where welfare rises by 12.92%, while the largest welfare decline for white workers also occurs in the middle quintile.
@@ 5 Business cycle dynamics
In this section, we first document the racial disparities over the business cycle. We then introduce aggregate uncertainty into our benchmark model to assess how hiring discrimination shapes racially disparate dynamics during economic fluctuations.
@@ 5.1 Racial disparities over the business cycles
Table 6 displays the labor market outcome disparities between black and white workers over the business cycle. The top panel presents the dynamics of unemployment rates for both groups, while the bottom panel shows the evolution of average wages after adjusting for composition bias. In this table, volatility is defined as the standard deviation of a variable relative to that of real GDP, and cyclicality is measured as the correlation between the variable and real GDP.
Unemployment rate: We use quarterly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the business cycle dynamics of unemployment rates. The statistics are computed using log deviations from an HP-filtered trend with a smoothing parameter of 1600. We find that both black and white unemployment rates are countercyclical, although the black unemployment rate is slightly less cyclical and less volatile than the white rate. These findings align with Nakajima (2021). In addition, the black-white unemployment rate gap is itself countercyclical: it widens during recessions and narrows during expansions, indicating that black workers see a larger unemployment increase in recessions but a faster decrease in expansions. This result is consistent with the countercyclical gap documented by Cajner et al. (2017).
@@
Black 9.51 -0.85
White 11.02 -0.92
Black-White gap 9.46 -0.63
@@
Black 2.54 2.60
White 2.08 1.27
Note: This table provides business cycle statistics for the US. The unemployment rate is obtained from the Current Population Survey (CPS) for 1996 to 2014. Average wage statistics are constructed from the Panel Study of Income Dynamics (PSID) for the same period. Unemployment rate statistics are calculated using HP filtered series with a smoothing parameter of 1600, while wage statistics are obtained using wage growth rates after controlling for demographics.
Wages: We use individual-level data from the PSID (1996–2014) to estimate the cyclical properties of black and white wages. Previous research (Stockman, 1983; Bils, 1985; Solon et al., 1994) shows that using aggregate data can be problematic due to changes in worker composition over the business cycle. Specifically, aggregate data tends to weight lower-skilled (and lower-wage) workers more heavily during expansions than recessions, causing countercyclical bias. To address this issue, we rely on the longitudinal structure of the PSID data to keep the composition of workers fixed over time. Following Bils (1985), Solon et al. (1994), and Devereux (2001), we estimate wage cyclicality by regressing wage growth on real GDP growth, controlling for observable characteristics. We measure wage volatility as the standard deviation of wage growth after controlling for demographics. [^13] Appendix D provides details on measuring the cyclicality and volatility of real wages. Our findings indicate that both black and white wages are procyclical, with black workers’ wages being twice as cyclical as those of white workers. Black workers’ wages also exhibit greater volatility than white workers’ wages.
@@ 5.2 Augmented model with aggregate shocks
We introduce shocks to aggregate total factor productivity (TFP) $z$ in our benchmark steady-state model described in Section 2. The TFP shocks follow an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$. The output of the matched worker-firm pairs depends on the realizations of both aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
With the introduction of aggregate shocks, the state space expands to include aggregate states $(z,\mu)$, where $\mu$ is the distribution of workers over their employment status ($e$), race ($R$), idiosyncratic productivity ($s$), and asset ($a$). The aggregate distribution for the next period $\mu^{\prime}$ is determined by $(z,\mu)$, and the endogenous law of motion given by $\mu^{\prime}=\Gamma(z,\mu)$ is determined in the equilibrium.
We follow Cooley et al. (1995) and Boppart et al. (2018) in choosing the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015. All other parameters remain the same as our benchmark calibration given in Table 1. We obtain the stochastic equilibrium of our model using the sequence space method proposed by Boppart et al. (2018). We first solve non-linearly for the perfect foresight transitions to a single small MIT shock, i.e., an unexpected shock to the aggregate TFP. We then use the solved impulse responses as the numerical derivatives with respect to the initial TFP shock. Using these derivatives, we simulate the stochastic equilibrium by generating the TFP realizations and calculating the corresponding model moments over the business cycles as a linear combination of the impulse response and the TFP realizations. [^14] This method hinges on the assumption that the business cycle dynamics can be well approximated as a linear system. In Appendix E, we demonstrate the validity of this assumption by establishing the symmetry of the impulse responses to 1% positive and negative TFP shocks.
@@
Black 9.51 0.26 -0.85 -0.69
White 11.02 0.22 -0.92 -0.71
Black-White gap 9.46 0.30 -0.63 -0.68
@@
Black 2.54 1.08 2.60 1.08
White 2.08 1.00 1.27 1.00
Note: This table compares the data moments with the moments generated by the benchmark model on unemployment and wages.
Table 7 compares business cycle statistics for unemployment and wages between the data and our model. Overall, the model replicates the key patterns observed in the data. It generates countercyclical unemployment rates, with white unemployment being slightly more cyclical than black unemployment, and it accurately reproduces the countercyclical gap between the two groups. In terms of wages, the model produces procyclical patterns, with black wages showing higher cyclicality and volatility than white wages. However, as is common with standard search models, our model falls short in generating realistic volatility levels for both black and white unemployment rates - a limitation discussed in Shimer (2005).
@@ 5.3 Impact of hiring discrimination over business cycle
We compare the benchmark dynamics with a counterfactual model that excludes prejudiced firms to examine how discriminatory hiring impacts the economy under aggregate fluctuations. In this one-sector counterfactual, only non-prejudiced firms operate, and workers face race-specific job separation rates as with Table 4.
@@ 5.3.1 Impulse responses
Figure 2 shows how labor market outcomes, wealth, and consumption for both black and white workers respond to an unexpected 1% expansionary TFP shock. We compare the benchmark model’s reaction with that of a counterfactual model that excludes $p$ firms. In both models, unemployment rates move countercyclically while wages, wealth, and consumption follow procyclical patterns. However, the responses in the counterfactual economy are generally less pronounced.
@@
Note: This figure plots the impulse response functions of unemployment rates, black-white average wage and wealth ratios, and aggregate consumption of black and white workers to a one percent increase in aggregate TFP at date zero. The red solid line denotes the responses from the benchmark model, and the blue dashed line denotes the responses from the counterfactual model without $p$ firms.
In the model without $p$ firms, the unemployment rate for black workers falls less after an expansionary shock than in the benchmark, indicating that black unemployment is less volatile when discriminatory hiring is removed. In the benchmark economy, $np$ firms compete with $p$ firms for hires, which leads them to post more vacancies and create additional opportunities for workers. Although white workers follow a similar pattern, the difference between the benchmark and counterfactual responses is much smaller for them. Notably, the counterfactual model shows a much faster recovery of unemployment rates than the benchmark model. [^15] This is a common feature in dual labor market models, where a single-market setup typically reverts to its steady state faster than a dual-market framework (e.g., Horvath and Yang, 2022).
The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than the model without prejudiced firms. This suggests more volatile procyclical wage and wealth responses for black workers than for white workers in the presence of $p$ firms. Moreover, consumption among both black and white workers responds more strongly in the benchmark model, leading to a more volatile consumption reaction following the expansionary TFP shock.
@@ 5.3.2 Stochastic simulation
We simulate the stochastic equilibrium of the economy over 10,000 periods under aggregate TFP shocks, and compare the outcomes with and without $p$ firms. While impulse responses capture the transition path following a one-time shock, stochastic equilibrium focuses on the economy’s short-term responses to a series of aggregate shocks. The results are summarized in Table 8.
The stochastic simulation results further support the insights from the impulse response analysis. We find that the presence of $p$ firms increases unemployment rate volatility for both black and white workers, but the effect is more pronounced for black workers. In the counterfactual economy without $p$ firms, both groups have an unemployment rate relative volatility of 0.18. With $p$ firms in the benchmark economy, the volatility for black workers rises to 0.26, compared to 0.22 for white workers. Similarly, the volatility of the black-white unemployment rate gap increases from 0.17 to 0.30, reflecting the disproportionate impact on black workers.
The wages of both black and white workers have dynamics similar to their unemployment rates. In particular, the presence of $p$ firms in the benchmark increases both the cyclicality and volatility of wages compared to the counterfactual without $p$ firms. However, this effect is more pronounced for black workers than for white workers. In our comparison of the benchmark and counterfactual economies, the inclusion of $p$ firms raises the cyclicality and volatility of black wages from 1.01 to 1.08, while white wages increase from 0.98 to 1. Overall, hiring discrimination not only heightens wage fluctuations but also widens the gap between black and white workers. These results support the findings of Cajner et al. (2017), Boulware and Kuttner (2019, 2024), Ganong et al. (2020), and Ragusett (2022), who argue that demographic differences alone cannot account for the racial disparities observed over the business cycle.
Table 8: Impact of hiring discrimination over business cycle
Volatility Cyclicality
@@
Black 0.26 0.18 -0.69 -0.70
White 0.22 0.18 -0.69 -0.70
Black-White gap 0.30 0.17 -0.67 -0.71
@@
Black 1.08 1.01 1.08 1.01
White 1.00 0.98 1.00 0.98
Average wealth
Black 0.39 0.38 0.12 0.12
White 0.24 0.24 0.07 0.07
Wealth ratio 0.16 0.14 0.21 0.22
Average consumption
Black 0.57 0.52 0.95 0.94
White 0.39 0.36 0.93 0.92
Consumption ratio 0.18 0.17 0.97 0.97
Note: This table compares the responses of black-white unemployment rates, wages, consumption, and wealth gaps over the business cycle between the benchmark model and the model without $p$ firms.
We further examine the cyclical properties of wealth and consumption for black and white workers. Consistent with our earlier findings on labor market outcomes, the presence of $p$ firms increases the volatility of both the wealth and consumption ratios. Specifically, with $p$ firms, the volatility of the wealth ratio rises from 0.14 to 0.16, and that of the consumption ratio increases from 0.17 to 0.18. These results indicate that, beyond labor market variables, the volatility of racial gaps in other dimensions is also notably reduced when discriminatory hiring is eliminated.
In summary, our analysis shows that hiring discrimination amplifies racial disparities over the business cycle. The presence of $p$ firms increases the volatility of outcomes for black workers relative to white workers across multiple dimensions, including labor market variables, wealth, and consumption. These findings suggest that discriminatory hiring plays a key role in driving the cyclical differences in economic outcomes between racial groups.